# Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review

### v1

@@ Abstract
Peer review is a critical process for ensuring the integrity of published scientific research. Confidence in this process is predicated on the assumption that experts in the relevant domain give careful consideration to the merits of manuscripts which are submitted for publication. With the recent rapid advancements in large language models (LLMs), a new risk to the peer review process is that negligent reviewers will rely on LLMs to perform the often time consuming process of reviewing a paper. However, there is a lack of existing resources for benchmarking the detectability of AI text in the domain of peer review. To address this deficiency, we introduce a comprehensive dataset containing a total of 788,984 AI-written peer reviews paired with corresponding human reviews, covering 8 years of papers submitted to each of two leading AI research conferences (ICLR and NeurIPS). We use this new resource to evaluate the ability of 18 existing AI text detection algorithms to distinguish between peer reviews written by humans and different state-of-the-art LLMs. Motivated by the shortcomings of existing methods, we propose a new detection approach which surpasses existing methods in the identification of AI written peer reviews. Our work reveals the difficulty of identifying AI-generated text at the individual peer review level, highlighting the urgent need for new tools and methods to detect this unethical use of generative AI.
Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review
Sungduk Yu Intel Labs sungduk.yu@intel.com                        Man Luo Intel Labs man.luo@intel.com                        Avinash Madusu Intel Labs avinash.madasu@intel.com
Vasudev Lal Intel Labs vasudev.lal@intel.com                        Phillip Howard Intel Labs phillip.r.howard@intel.com
@@ 1 Introduction
Figure 1: Left panel shows the data construction pipeline and the right panel shows the proposed method Anchor Embedding Approach for detecting AI generated review.
@@
Recent advancements in large language models (LLMs) have enabled their application to a broad range of domains, where LLMs have demonstrated the ability to produce plausible and authoritative responses to queries even in highly technical subject areas. These advancements have coincided with a surge in interest in AI research, resulting in increased paper submissions to leading AI conferences (Audibert et al., 2022). Consequently, workloads for peer reviewers have also increased significantly, which could make LLMs an appealing tool for lessening the burden of fulfilling their peer review obligations (Kuznetsov et al., 2024; Kousha and Thelwall, 2024; Zhuang et al., 2025).
@@
Despite their impressive capabilities, the use of LLMs in the peer review process raises several ethical and methodological concerns which could compromise the integrity of the publication process (Hosseini and Horbach, 2023; Latona et al., 2024; Seghier, 2024; Zhou et al., 2024). Reviewers are selected based on their expertise in a technical domain related to a submitted manuscript, which is necessary to critically evaluate the proposed research. Offloading this responsibility to an LLM circumvents the role that reviewer selection plays in ensuring proper vetting of a manuscript. Furthermore, LLMs are prone to hallucination and may not possess the ability to rigorously evaluate research publications. Therefore, the use of LLMs in an undisclosed manner in peer review poses a significant ethical concern that could undermine confidence in this important process.
@@
Motivating the need for evaluation resources and detection tools to address this problem is the apparent increase in AI-generated text among peer reviews submitted to recent AI research conferences. Recent studies revealed an upward trend in AI-generated texts among peer reviews at the corpus level (Liang et al., 2024a; Latona et al., 2024). This trend is particularly concerning given that evaluations from human and AI reviewers are not aligned (Drori and Te’eni, 2024; Latona et al., 2024; Ye et al., 2024), suggesting that the unregulated and undisclosed use of LLMs in peer review could undermine the integrity of the current system.
@@
Despite the growing recognition of this problem, there is a lack of existing dataset resources for comprehensively evaluating the performance of AI text detection methods in the domain of peer review. To address this deficiency, we introduce the largest dataset to-date of parallel human-written and LLM-written peer reviews for identical papers submitted to 8 years of papers for two leading AI research conferences, NeurIPS and ICLR (Figure 1). Our dataset consolidates human-written peer reviews from existing sources with AI-written peer reviews that we generated for the same paper using five state-of-the-art LLMs: GPT-4o Achiam et al. (2023), Claude Sonnet 3.5 Anthropic (2023), Gemini 1.5 pro Team et al. (2023), Qwen 2.5 72b Bai et al. (2023), and Llama 3.1 70b Dubey et al. (2024). In total, our dataset contains 788,984 peer reviews, evenly balanced between human-written reviews and AI-generated peer reviews created by these five LLMs.
@@
We use our dataset to investigate the suitability of various AI text detection methods for identifying LLM generations in the peer review process. While limited prior work has analyzed the presence of AI-generated text in peer reviews at the corpus level (Liang et al., 2024a) or used conference-submitted reviews only (Latona et al., 2024), our study is the first to investigate the detectability LLM generations at the individual review level using systematically generated AI samples along with human samples, which is necessary to address this problem in practice. Specifically, we evaluate 18 open-source methods.
@@
Our results show that most existing AI-text detection methods are limited in their ability to robustly detect AI-generated reviews while maintaining a low number of false positives. Motivated by this finding, we propose an alternative approach to detecting AI-generated peer reviews by comparing the semantic similarity of a given review to a set of reference AI-generated reviews for the same paper, which surpasses the performance of all existing approaches in detecting GPT-4o and Claude written peer reviews. We conduct analyses to understand how different levels of AI use for editing reviews impacts detectability and false positives, as well as the characteristics which distinguish LLM-written peer reviews from those written by humans. Our work demonstrate the challenge of detecting AI-written text in peer reviews and motivates the need for further research on methods to address this unethical use of LLMs in the peer review process.
@@
To summarize, our contributions are as follows: (1) We introduce a new dataset containing 788,984 AI-written peer reviews generated by five widely-used LLMs along with corresponding human-written peer reviews for the same papers, which is the largest resource of its kind for evaluating the detectability of AI text in peer review. (2) Using our dataset, we evaluate 18 existing open source AI text detection algorithms as well as a commercial API, finding that most methods struggle to reliably detect fully AI-written peer reviews at low levels of false positives. (3) We propose a new detection method which compares the semantic similarity of a given peer review to a reference LLM-generated peer review for the same paper, which outperforms all existing detection approaches in our evaluations. (4) We conduct analyses to highlight the characteristics which distinguish human-written peer reviews from those generated by LLMs, show LLM reviews are generally more favorable and more confident it their assessment, and evaluate how the use of LLMs for editing peer reviews impact false positive rates. (5) We make our dataset publicly available under an open source license to facilitate future work on detecting AI generated text in peer review.
@@ 2 Related Work
Several benchmark datasets have been introduced to evaluate AI-based text detection models. RAID-TD  Dugan et al. (2024) provides a large-scale benchmark designed to assess text detection under adversarial conditions, ensuring robustness against manipulated AI-generated content. The M4 Dataset  Wang et al. (2024) expands the scope by incorporating reviews from multiple LLMs across different languages, offering a more diverse linguistic evaluation. The HC3 Dataset  Guo et al. (2023) consists of responses from ChatGPT and human experts, covering specialized domains such as finance, medicine, and law, in addition to general open-domain content. In contrast, the GPT Reddit Dataset (GRiD)  Qazi et al. (2024) focuses on social media conversations, compiling a diverse set of human- and AI-generated responses to Reddit discussions. Meanwhile, Beemo Artemova et al. (2024) introduces a benchmark of expert-edited machine-generated outputs, spanning creative writing, summarization, and other practical applications, further refining detection challenges. These benchmarks primarily evaluate AI-generated text from a single model and do not address the domain of AI text in peer review. In contrast, our dataset is larger than most existing datasets (788k generations) and is unique in its focus on AI text detection in peer review.
@@
AI-generated text detection has been framed as a binary classification task to distinguish human-written from machine-generated text  Bakhtin et al. (2019); Jawahar et al. (2020); Fagni et al. (2021); Mitchell et al. (2023a). Solaiman et al. (2019) used a bag-of-words model with logistic regression for GPT-2 detection, while fine-tuned language models like RoBERTa Liu et al. (1907) improved accuracy  Zellers et al. (2019); Uchendu et al. (2020); Gehrmann et al. (2019). Zero-shot methods based on perplexity and entropy emerged as alternatives  Ippolito et al. (2020); Gehrmann et al. (2019). Other studies focused on linguistic patterns and syntactic features for model-agnostic detection  Uchendu et al. (2020); Gehrmann et al. (2019). Watermarking techniques, such as DetectGPT  Mitchell et al. (2023a), have also been proposed for proactive identification. Centralized frameworks like MGTBench  He et al. (2023) and its refined version, IMGTB  Spiegel and Macko (2023), provide standardized evaluations for AI text detection. IMGTB categorizes methods into model-based and metric-based approaches. Model-based methods leverage large language models such as ChatGPT-turbo  OpenAI (2023) and Claude  Anthropic (2023). Metric-based methods, including Log-Likelihood  Solaiman et al. (2019), Rank  Gehrmann et al. (2019), Entropy  Gehrmann et al. (2019), DetectGPT  Mitchell et al. (2023a), and LRR  Su et al. , rely on log-likelihood and ranking for classification.
@@
Recent studies have explored the role of LLMs in peer review, examining their influence on reviewing practices  Liang et al. (2024a, b), simulating multi-turn interactions  Tan et al. (2024), and assessing their reviewing capabilities  Zhou et al. (2024). Tools like OpenReviewer  Tyser et al. provide AI-assisted review improvements, while other works focus on LLM transparency  Kuznetsov et al. (2024) and distinguishing AI-generated content  Mosca et al. (2023). Other recent work has investigated AI-driven review systems  Tyser et al. (2024), agentic frameworks  Liang et al. (2024b), and comparative analyses of LLM-generated reviews  Liu and Shah (2023), along with broader explorations of LLMs’ role in peer review  Jin et al. (2024); Santu et al. (2024); Sukpanichnant et al. (2024). While it is not the primary focus of our work, we analyze the quality of LLM-generated peer reviews in Section 6.
@@
3 Methods and Dataset Construction
3.1 Human reviews
We used the OpenReview API (OpenReview, ) to collect submitted manuscripts and their reviews for the ICLR conferences from 2019 to 2024, as well as for NeurIPS conferences from 2021 to 2024. Additionally, we used the ASAP dataset Yuan et al. (2022) to collect manuscripts and reviews for ICLR 2017 to 2018 and NeurIPS 2016 to 2019111NeurIPS 2020 reviews are not publicly available.
3.2 AI reviews
Prompts. To control the content and structure of these AI-generated reviews, we included conference-specific reviewer guidelines and review templates in the prompts. Note that this is nontrivial since the review templates have evolved significantly over the years (Table 5). In addition, we aligned the paper decisions by prompting the LLMs with specific decisions derived from the corresponding human reviews (see Appendix E for complete prompt details). This step is important, as we found that AI review content and recommendations vary substantially depending on how input prompts are constructed. Thus, these measures represent our efforts to control the influence of text prompts.
Computation. We used Azure OpenAI Service, Amazon Web Services, and Google Cloud Platform to generate GPT-4o, Claude, and Gemini reviews, respectively. For Qwen reviews, we used NVIDIA RTX 6000 GPUs, and for Llama reviews, we used Intel Gaudi-2 accelerators.
3.3 Dataset Statistics
@@
We note slight variations in sample sizes among the five LLMs. These differences stem from LLM-specific factors such as context window limits, generation errors (e.g., malformed JSON or gibberish outputs), and input prompt safety filters.
3.4 Review Editing
LLMs have been widely used for editing human writing Laban et al. (2024); Raheja et al. (2024), including tasks such as grammar checking and enhancing language fluency, which are particularly beneficial for non-native speakers in conveying their reviews more clearly. However, some users may overly rely on LLMs, leading to significant alterations of the original review. To simulate such scenarios in our dataset, we generate edited reviews with four varying levels of modification using LLMs (minimum, moderate, extensive, and maximum). This approach enables us to assess how robust detection methods are for text with different degrees of editing. Appendix E.3 illustrates the types of prompts used to generate these edits.
To validate whether these prompts produce reviews with varying levels of modification, we conducted a similarity checking experiment. Our hypothesis is that the similarity between the edited review and the original review decreases as the level of modification increases, ranging from minimal edits to extensive revisions. To test this, we used a sentence embedding model to compute embeddings for both the original and edited reviews and calculated their average cosine similarity scores. The results, summarized in Table 2, confirm that the editing levels are indeed distinct.
Minimum Moderate Extensive Maximum
0.9841 0.9261 0.8616 0.6799
Table 2: Similarity score of edited reviews and original human-written reviews using different prompts.
4 Anchor Embeddings Detection Methodology
We propose a novel method for detecting AI-generated reviews by comparing their semantic similarity to a reference AI-generated review for the same article, referred to as the “Anchor Review" (AR). The AR can be generated by any LLM. We use a simple, generic prompt (Appendix E.2) to generate the AR without prior knowledge of the user prompts (i.e., the AR prompt differs from those used to create reviews in the testing dataset). Once an AR is generated for a given paper (Eq.1), and a testing review (TR) is provided, we obtain their embeddings using a text embedding model (EM, Eqs.2 and 3). The semantic similarity between the embeddings of the AR and TR is then computed using the cosine similarity function (Eq.4). Finally, this similarity score is compared against a learned threshold ($\theta$): if the score exceeds the threshold, the review is classified as AI-generated; otherwise, it is not (Eq. 5). Putting everything together, the method is formalized as:
$\text{AR}=\text{LLM}(\text{paper},\text{Prompt}_{\text{AR}})$ (1)
[…]
In our study, we use OpenAI’s embedding model (text-embedding-003-small). The threshold $\theta$ is learned from the calibration set. Specifically, for each review in the calibration data, we apply the steps outlined in Eqs. 1 to 4.
Voting of Multiple Anchors
Intuitively, our approach performs best when the anchor embeddings are generated using the same model that produced the test review (source LLM). However, in real-world scenarios, the source LLM is typically unknown (a situation commonly referred to as “black-box” detection scenario). To address this challenge, we propose a voting-based technique. Specifically, we generate multiple anchor embeddings using different types of LLMs (anchor LLMs). For each anchor embedding, we compute the Score (Eq.4) and derive the corresponding label assignment (Eq.5). If at least one anchor embedding assigns a positive label, the final label is positive. Otherwise, the final label is negative.
In our experiments in the following sections, we use three anchor reviews for voting—each generated by GPT-4o, Gemini, and Claude, respectively.
@@ 5.1 Fully AI-Written Review Detectability
We compare our approach with 18 baseline methods from IGMBT (see Appendix A for details) and utilize the calibration set to determine appropriate thresholds for the test data. The threshold is then determined by setting an target False Positive Rate (FPR), which is achieved by adjusting the threshold until the FPR equals the target value. In our setup, we have studied varying thresholds (e.g., 0.1%, 0.5%, and 1%). We focus on low FPR targets because false positive classifications—where human-written reviews are mistakenly identified as AI-generated—carry high stakes, potentially damaging an individual’s reputation. Besides, we focus on AI review text samples generated by three commercial LLMs (GPT-4o, Gemini, and Claude) because these models are more advanced, making the text detection task harder. Also, general users are more likely to choose them over open-source LLMs due to their convenient user interfaces and limited access to the compute resources required for running open-source models.
@@
AI text detection models can be calibrated for varying levels of sensitivity in order to balance the trade-off between true positive and false positive detections. Receiver operating characteristic (ROC) curves are therefore commonly used to compare different methods, as they provide a visualization of the true-positive rate (TPR) which can be achieved by a model when its decision boundary is calibrated for a range of different false positive rates (FPR). Figure 2 provides the ROC curves for our anchor embedding approach and baseline methods, calculated using our GPT-4o, Gemini, and Claude review calibration subset separately for reviews submitted to ICLR2021-2022 (left) and NeurIPS2021-2022 (right). The area under the curve (AUC) is provided for each method in the legend; higher values indicate better detection accuracy across the entire range of FPR values. These plots show that our anchor embedding approach consistently achieves the highest AUC for AI peer review detection among evaluated methods, followed by the baseline method Binoculars.
@@
While ROC curves are useful for comparing the overall performance of different classifiers, only the far left portion of these plots are typically relevant for practical applications of AI text detection models. This is particularly true in the domain of peer review, where the cost of a false positive is high. Reviewers volunteer their time and expertise to sustain this important process; false accusations have the potential to further reduce the availability of reviewers due to disengagement and can also lead to significant reputational harm. Therefore, it is vital that AI text detection systems for peer review be calibrated for a low FPR in order to avoid such negative outcomes.
@@
Prior work has shown that AUC is not necessarily reflective of how models perform at very low FPR values (Yang et al., 2023; Krishna et al., 2024; Tufts et al., ). Therefore, we also report the actual TPR and FPR achieved by different detection methods at discrete low values of target FPR (0.1%, 0.5%, and 1%), which we believe to be of greatest interest for practical applications. The target FPR is used to calculate each method’s classification threshold using our calibration dataset, with the actual TPR and FPR computed over the withheld test dataset. To simulate a more challenging evaluation setting where some of the test reviews are “out-of-domain” in the sense that they come from a different conference than the calibration dataset, we use only the ICLR reviews to calibrate each method (see Section C.1 for in-domain evaluations). Table 3 provides these results separately for the detection of GPT-4o, Gemini, and Claude reviews (Llama and Qwen results are provided in Appendix C.2). Other baseline methods that failed achieve a TPR of at least 1% at a target FPR of 1% (i.e., TPR@1%FPR < 1%) are omitted.
The results in Table 3 show that our anchor embedding approach achieves the highest TPR at low FPR values among evaluated methods for GPT-4o and Claude-generated reviews. This is notable, as GPT-4o and Claude are among the most advanced and widely-used LLMs currently available. The performance difference between our approach and the next best-performing method (Binoculars) is particularly large for GPT-4o reviews and low target FPR settings (e.g., 0.1%), with our anchor embedding approach achieving absolute improvements of up to 50.1% in TPR. Furthermore, our anchor embedding approach exactly achieves the target FPR on the test set for these reviews, whereas other methods exceed the target FPR by as much as 0.7%. Gemini-generated reviews appear to be easier for baseline methods to detect than GPT-4o and Claude reviews, and we observe that our anchor embedding approach performs similarly on these reviews as Binoculars. This suggests that the anchor embedding method provides the greatest benefits over existing methods for reviews which are the most challenging to detect; in fact, our approach performs best for detecting GPT-4o reviews, whereas other strong baselines such as Binoculars perform the worst on these reviews.

### v2

# Is Your Paper Being Reviewed by an LLM? Benchmarking AI Text Detection in Peer Review

@@ Abstract
Peer review is a critical process for ensuring the integrity of published scientific research. Confidence in this process is predicated on the assumption that experts in the relevant domain give careful consideration to the merits of manuscripts which are submitted for publication. With the recent rapid advancements in large language models (LLMs), a new risk to the peer review process is that negligent reviewers will rely on LLMs to perform the often time consuming process of reviewing a paper. However, there is a lack of existing resources for benchmarking the detectability of AI text in the domain of peer review. To address this deficiency, we introduce a comprehensive dataset containing a total of 788,984 AI-written peer reviews paired with corresponding human reviews, covering 8 years of papers submitted to each of two leading AI research conferences (ICLR and NeurIPS). We use this new resource to evaluate the ability of 18 existing AI text detection algorithms to distinguish between peer reviews fully written by humans and different state-of-the-art LLMs. Additionally, we explore a context-aware detection method called Anchor, which leverages manuscript content to detect AI-generated reviews, and analyze the sensitivity of detection models to LLM-assisted editing of human-written text. Our work reveals the difficulty of identifying AI-generated text at the individual peer review level, highlighting the urgent need for new tools and methods to detect this unethical use of generative AI. Our dataset is publicly available at: https://huggingface.co/datasets/IntelLabs/AI-Peer-Review-Detection-Benchmark.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled their application to a broad range of domains, where LLMs have demonstrated the ability to produce plausible and authoritative responses to queries even in highly technical subject areas. These advancements have coincided with a surge in interest in AI research, resulting in increased paper submissions to leading AI conferences [1]. Consequently, workloads for peer reviewers have also increased significantly, which could make LLMs an appealing tool for lessening the burden of fulfilling their peer review obligations [2, 3, 4].
Figure 1: Left panel: our data construction pipeline. Right panel: our context-aware detection method (Anchor) specifically designed for AI-generated review detection evaluated in Section 5.1.
@@
Despite their impressive capabilities, the use of LLMs in the peer review process raises several ethical and methodological concerns which could compromise the integrity of the publication process [5, 6, 7, 8]. Reviewers are selected based on their expertise in a technical domain related to a submitted manuscript, which is necessary to critically evaluate the proposed research. Offloading this responsibility to an LLM circumvents the role that reviewer selection plays in ensuring proper vetting of a manuscript. Furthermore, LLMs are prone to hallucination and may not possess the ability to rigorously evaluate research publications. Therefore, the use of LLMs in an undisclosed manner in peer review poses a significant ethical concern that could undermine confidence in this important process.
@@
Motivating the need for evaluation resources and detection tools to address this problem is the apparent increase in AI-generated text among peer reviews submitted to recent AI research conferences. Prior studies revealed an upward trend in AI-generated texts among peer reviews [9, 6]. This trend is particularly concerning given that evaluations from human and AI reviewers are not aligned [10, 6, 11] and that LLM-generated reviews lack robustness [12], suggesting the unregulated and undisclosed use of LLMs in peer review could undermine the integrity of the current system.
@@
Despite the growing recognition of this problem, there is a lack of existing dataset resources for comprehensively evaluating the performance of AI text detection methods in the domain of peer review. To address this deficiency, we introduce the largest dataset to-date of parallel human-written and LLM-written peer reviews for 8 years of papers submitted to two leading AI research conferences, NeurIPS and ICLR (Figure 1). Our dataset consolidates human-written peer reviews from existing sources with AI-written peer reviews that we generated for the same paper using five state-of-the-art LLMs: GPT-4o [13], Claude Sonnet 3.5 [14], Gemini 1.5 pro [15], Qwen 2.5 72b [16], and Llama 3.1 70b [17]. In total, our dataset contains 788,984 peer reviews, evenly balanced between human-written reviews and AI-generated peer reviews created by these five LLMs.
@@
We use our dataset to investigate the suitability of various AI text detection methods for identifying LLM generations in the peer review process. While limited prior work has analyzed the presence of AI-generated text in peer reviews at the corpus level [9] or has analyzed the use of propriety solutions [6], our study is the first to investigate the detectability LLM generations at the individual review level using synthetically generated AI samples, which is necessary to address this problem in practice. Specifically, we evaluate 18 existing open-source methods for AI text detection.
@@
Our results show that most existing AI-text detection methods are limited in their ability to robustly detect AI-generated reviews while maintaining a low number of false positives. Motivated by this finding, we test an alternative approach which is specifically designed for AI text detection in the peer review context. Leveraging the additional context available in the peer review setting, our method detects AI-generated peer reviews by comparing the semantic similarity of a given review to a set of reference AI-generated reviews for the same paper. We find that this simple yet effective method surpasses the performance of all existing approaches in detecting GPT-4o and Claude written peer reviews. Additionally, we conduct analyses to understand how different levels of AI use for editing reviews impacts detectability and false positives, as well as the characteristics which distinguish LLM-written peer reviews from those written by humans. Our work demonstrate the challenge of detecting AI-written text in peer reviews and motivates the need for further research on methods to address this unethical use of LLMs in the peer review process.
@@
To summarize, our contributions are as follows: (1) We publicly release a dataset of 788,984 AI-written peer reviews generated by five widely-used LLMs paired with human-written reviews for the same papers, which is the largest resource to-date for studying AI text detection in peer review. (2) Using our dataset, we benchmark 18 open-source AI text detection algorithms, finding that most struggle to reliably detect fully AI-written peer reviews at low false positive rates. (3) We propose a new context-aware detection method which compares the semantic similarity between a candidate review and a reference LLM-generated review for the same paper, achieving strong performance under strict FPR constraints. (4) We conduct analyses revealing key differences between human- and AI-written reviews, finding that AI-generated reviews are generally less specific, more favorable, and more confident. (5) We evaluate how LLM-assisted editing in peer review affects detection rates.
@@ 2 Related Work
Several datasets have been introduced to evaluate AI text detection models. RAID-TD  [18] provides a large-scale benchmark designed to assess text detection under adversarial conditions, ensuring robustness against manipulated AI-generated content. The M4 Dataset  [19] expands the scope by incorporating reviews from multiple LLMs across different languages, offering a more diverse linguistic evaluation. The HC3 Dataset  [20] consists of responses from ChatGPT and human experts, covering specialized domains such as finance, medicine, and law, in addition to general open-domain content. In contrast, the GPT Reddit Dataset (GRiD)  [21] focuses on social media conversations, compiling a diverse set of human- and AI-generated responses to Reddit discussions. Meanwhile, Beemo [22] introduces a benchmark of expert-edited machine-generated outputs, spanning creative writing, summarization, and other practical applications. These benchmarks primarily evaluate AI-generated text from a single model and do not address the domain of AI text in peer review. In contrast, our dataset is larger than most existing datasets (788k generations) and is unique in its focus on AI text detection in peer review.
@@
AI-generated text detection has been framed as a binary classification task to distinguish human-written from machine-generated text  [23, 24, 25, 26]. Solaiman et al. [27] used a bag-of-words model with logistic regression for GPT-2 detection, while fine-tuned language models like RoBERTa [28] improved accuracy  [29, 30, 31]. Zero-shot methods based on perplexity and entropy emerged as alternatives  [32, 31]. Other studies focused on linguistic patterns and syntactic features for model-agnostic detection  [30, 31]. Watermarking techniques, such as DetectGPT  [26], have also been proposed for proactive identification. Centralized frameworks like MGTBench  [33] and its refined version, IMGTB  [34], provide standardized evaluations for AI text detection. IMGTB categorizes methods into model-based and metric-based approaches. Model-based methods leverage large language models such as ChatGPT-turbo  [35] and Claude  [14]. Metric-based methods, including Log-Likelihood  [27], Rank  [31], Entropy  [31], DetectGPT  [26], and LRR  [36], rely on log-likelihood and ranking for classification.
@@
Recent studies have explored the role of LLMs in peer review, examining their influence on reviewing practices  [9, 37], simulating multi-turn interactions  [38], and assessing their reviewing capabilities  [8]. Tools like OpenReviewer  [39] provide AI-assisted review improvements, while other works focus on LLM transparency  [2] and distinguishing AI-generated content  [40]. Recent studies have investigated AI-driven review systems  [41], agentic frameworks  [42], and comparative analyses of LLM-generated reviews  [43], along with broader explorations of LLMs’ roles and limitations in peer review  [44, 45, 46, 37]. While it is not the primary focus of our work, we analyze the quality of LLM-generated peer reviews in Section 5.
3 Dataset Construction
3.1 Human reviews
We used the OpenReview API [47] to collect submitted manuscripts and their reviews for the ICLR conferences from 2019 to 2024, as well as for NeurIPS conferences from 2021 to 2024111NeurIPS 2020 reviews are not publicly available. Additionally, we used the ASAP dataset [48] to collect manuscripts and reviews for ICLR 2017 to 2018 and NeurIPS 2016 to 2019.
3.2 AI reviews
Prompts. To control the content and structure of these AI-generated reviews, we included conference-specific reviewer guidelines and review templates in the prompts. Note that review templates have evolved significantly over time (Table S1), necessitating prompt adaptations for papers submitted in different years. We additionally aligned the paper decisions by prompting the LLMs with specific decisions derived from the corresponding human reviews (see Appendix E for complete prompt details). This step is important, as we found that AI review content and recommendations vary substantially depending on how input prompts are constructed. Thus, these measures represent our efforts to control the influence of text prompts. Importantly, as we show later in Section 5.2, despite using a consistent prompting strategy, our dataset remains robust to prompt variation and supports generalizable detection performance across different prompting styles.
Computation. We used Azure OpenAI Service, Amazon Web Services, and Google Cloud Platform to generate GPT-4o, Claude, and Gemini reviews, respectively. For Qwen reviews, we used NVIDIA RTX 6000 GPUs, and for Llama reviews, we used Intel Gaudi 2 accelerators.
3.3 Dataset Statistics
@@
While we include reviews from post-ChatGPT [^2] ChatGPT was released on November 30, 2022. conferences (ICLR 2023–2024 and NeurIPS 2023–2024), we note that the human-labeled reviews from these years may include LLM-assisted texts. These reviews are retained to support broader research use cases, such as longitudinal analysis of linguistic trends. However, we do not use them in our main experiments, which focus on pre-2023 data where we have higher confidence in the human-authored labels.
We note slight variations in sample sizes among the five LLMs. These differences stem from LLM-specific factors such as context window limits, generation errors (e.g., malformed JSON or degenerate outputs), and input prompt safety filters.
@@ 4.1 Fully AI-Written Review Detectability
We compare our approach with 18 baseline methods using IMGTB (see Appendix A for details) and utilize the calibration set to determine appropriate thresholds for the test data. To minimize the risk of LLM-text contamination in human-written reviews, we only include reviews submitted before the release of ChatGPT in late 2022 (i.e., NeurIPS 2016-2022 and ICLR 2017-2022 papers), when LLM use rapidly became widespread. The threshold is then determined by setting a target False Positive Rate (FPR), which is achieved by adjusting the threshold until the FPR equals the target value. We focus on low FPR targets (e.g., 0.1%, 0.5%, and 1%) because false positive classifications—where human-written reviews are mistakenly identified as AI-generated—carry high stakes, potentially damaging an individual’s reputation. Additionally, we focus on AI review text samples generated by three commercial LLMs (GPT-4o, Gemini, and Claude) because these models are more advanced, making the AI text detection task harder. Many users are more likely to choose these models over open-source LLMs due to their convenient user interfaces and limited access to the compute resources required for running advanced open-source models.
@@
AI text detection models can be calibrated for varying levels of sensitivity in order to balance the trade-off between true positive and false positive detections. Receiver operating characteristic (ROC) curves are therefore commonly used to compare different methods, as they provide a visualization of the true-positive rate (TPR) which can be achieved by a model when its decision boundary is calibrated for a range of different FPRs. Figure 2 provides the ROC curves for baseline methods, calculated using our GPT-4o, Gemini, and Claude review calibration subset separately for reviews submitted to ICLR 2021-2022 (left) and NeurIPS 2021-2022 (right). The area under the curve (AUC) is provided for each method in the legend; higher values indicate better detection accuracy across the entire range of FPR values. Among the 18 baseline models, Binoculars consistently performs the best across the ROC curve, particularly excelling at maintaining a low FPR. Our task-specific anchor method appears to perform strongly, but we will discuss its performance separately in Section 5.1; here, we focus on analyzing the 18 existing detection methods.
@@
Although ROC curves are useful for comparing the overall performance of different classifiers, only the far left portion of these plots are typically relevant for practical applications of AI text detection models. This is particularly true in the domain of peer review, where the cost of a false positive is high. Reviewers volunteer their time and expertise to sustain this important process; false accusations have the potential to further reduce the availability of reviewers due to disengagement and can also lead to significant reputational harm. Therefore, it is vital that AI text detection systems for peer review be calibrated for a low FPR in order to avoid such negative outcomes.
@@
Prior work has shown that AUC is not necessarily reflective of how models perform at very low FPR values [49, 50, 51]. Therefore, we also report the actual TPR and FPR achieved by different detection methods at discrete low values of target FPR (0.1%, 0.5%, and 1%), which we believe to be of greatest interest for practical applications. The target FPR is used to calculate each method’s classification threshold using our calibration dataset, with the actual TPR and FPR computed over the withheld test dataset. To simulate a more challenging evaluation setting where some of the test reviews are “out-of-domain” in the sense that they come from a different conference than the calibration dataset, we use only the ICLR reviews to calibrate each method (see Section C.2 for in-domain evaluations).
Table 2: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Highest TPRs are in bold. The results are well separated; see Table S5 for uncertainties estimated via bootstrapping.